OpenStudy (anonymous):

Solve the equation. First, express your answer in terms of natural logarithms (for instance, x = (2 + ln(5)) / (ln(3)). 28e^-x/3=54.5 x= Now use a calculator to find an approximation for your answer. (Round your answer to two decimal places.) x =

OpenStudy (anonymous):

divide both sides by 28 then take ln to each side.

OpenStudy (anonymous):

-x/3 = ln (54.5/28) -x = 3ln(54.5/28)=ln{(54.5/28)}^3

OpenStudy (anonymous):

what about the e?

OpenStudy (anonymous):

multiply both sides by -1 and use the fact that -1lna = ln(a^-1)

OpenStudy (anonymous):

ln e = 1

OpenStudy (anonymous):

ln(e^-x/3)=(-x/3)lne=-x/3

OpenStudy (anonymous):

why dont we divide by -1?

OpenStudy (anonymous):

mutliply or divide both sides by -1 to solve for x

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (anonymous):

x = ln(28^3) - ln (54.5^3)

OpenStudy (anonymous):

or x = 3ln28 - 3ln54.5

OpenStudy (anonymous):

what about the second answer?

OpenStudy (anonymous):

they are both the same answer since ln(28^3) = 3ln28 and ln(54.5^3) = 3ln54.5

OpenStudy (anonymous):

i meant the answer for part 2 of the question

OpenStudy (anonymous):

type 3ln28 - 3ln54.5 into your calculator

OpenStudy (anonymous):

i dont see ln on my calc

OpenStudy (anonymous):

nevermind. thanks for all the help!